Find Two consecutive odd integars if twice the larger, increased by the smaller equals 85
Answer by Maths68(1474)

Let x+1 is smaller odd integer
and x+3 is larger odd integer
2(x+3)+(x+1)=85
2x+6+x+1=85
3x+7=85
3x=85-7
x=78/3
x=26
Smaller odd integer = x+1=26+1=27
Larger odd integer = x+3=26+3=29
27 and 29 are two consecutive odd integers